# AdaptNC: Adaptive Nonconformity Scores for Uncertainty-Aware Autonomous Systems in Dynamic Environments

### v1

@@ B.3 Dynamically Tuned Adaptive Conformal Inference (DtACI)
B.4 High Density Region Identification
@@
We analyze the convergence properties of the MCKDE method. Let the true underlying density function be denoted as $f$ on $\mathbb{R}^{d}$. For a specified miscoverage rate $\alpha\in(0,1)$, the true High-Density Region (HDR) is defined as the level set $R_{b}=\{z\in\mathbb{R}^{d}:f(z)\geq b\}$. We aim to show that the estimated region $\hat{R}_{N,M}$ converges to the true region $R_{\alpha}$ as both the number of samples $N$ used to construct the KDE and the number of Monte Carlo samples $M$ increase.
The procedure described in Algorithm 4 aims to estimate this region using a Kernel Density Estimator (KDE) $\hat{f}_{N}$ constructed from $N$ reweighted samples drawn from the true distribution. This reweighted empirical distribution is considered to be the ”true” distribution for the purpose of the monte-carlo sampling.
From this distribution, we draw $M$ Monte Carlo samples $\{z_{1},\ldots,z_{M}\}$. These are then scored using the KDE to obtain density estimates $\xi_{i}=\hat{f}_{N}(z_{i})$. The threshold $\hat{\tau}$ is determined as the $\alpha$-quantile of these scores, $\hat{\tau}=\textrm{Quantile}(\{\xi_{i}\}_{1\leq i\leq M},\alpha)$, and the estimated HDR is defined as $\hat{R}_{N,M}=\{z_{i}:\xi_{i}\geq\tau\}$.
Theorem B.1 (Consistency of MCKDE).
Given the following conditions:
  1. 1.
    The density $f$ is bounded and twice continuously differentiable.
  2. 2.
    The gradient of the density has strictly positive magnitude, $\inf_{f^{-1}(\{t\})}\|\nabla f\|>0$ for almost every $t\geq 0$.
  3. 3.
    The measure $\lambda(f^{-1}([t_{\alpha}-\epsilon,t_{\alpha}+\epsilon]))\to 0$ as $\epsilon\to 0$, and $\lambda(f^{-1}(0,\epsilon])\to 0$ as $\epsilon\to 0$.
  4. 4.
    The kernel function $K$ is the gaussian kernel.
  5. 5.
    The bandwidth of the KDE is selected according to the rules of Scott (1979) or Silverman (1986).
Then, as $N\to\infty$ and $M\to\infty$, the estimated region $\hat{R}_{N,M}$ converges to the true region $R_{\alpha}$ in probability:
$\lambda(\hat{R}_{N,M}\Delta R_{\alpha})\xrightarrow{P}0$ (9)
Proof.
We note that the procedure described in Algorithm 4 can be viewed as a two-step estimation process. First, we estimate the density function $f$ using a Kernel Density Estimator (KDE) $\hat{f}_{N}$ based on $N$ reweighted samples. Second, we use Monte Carlo sampling to estimate the level set of the KDE, $\hat{R}_{N,M}$. This two-step process allows us to analyze the convergence properties of each step separately.
Step 1: Convergence of the KDE Under the assumptions stated in Theorem B.1, particularly the smoothness of $f$ and the choice of bandwidth, Theorem B.3 shows that the KDE $\hat{f}_{N}$ converges to the true density $f$ almost surely as $N\to\infty$ . This implies the following:
$\lim_{N\to\infty}\|\hat{f}_{N}-f\|=0$ (10)
Therefore, as $N$ goes to $\infty$, the estimated density $\hat{f}_{N}$ is the same as the true density $f$.
We must also show that the level sets of $\hat{f}_{N}$ converge to those of $f$. Let $R_{\alpha}=\{x:f(x)\geq t_{\alpha}\}$ be the true high-density region, where $t_{\alpha}$ is the unique threshold such that $\mathbb{P}_{Z\sim f}(Z\in R_{\alpha})=1-\alpha$. Let $R_{N}=\{x\in\mathbb{R}^{d}:\hat{f}_{N}(x)\geq t_{N}\}$ denote the level set estimate, where $t_{N}$ is the unique threshold such that the region contains exactly $1-\alpha$ of the probability mass defined by $\hat{f}_{N}$:
$\mathbb{P}_{Z\sim\hat{f}_{N}}(Z\in R_{N})=\int_{\{z:\hat{f}_{N}(z)\geq t_{N}\}}\hat{f}_{N}(z)dz=1-\alpha$ (11)
By Corollary B.5, we have that the measure of the symmetric difference between the level sets converges in probability:
$\lambda(R_{N}\Delta R_{\alpha})\xrightarrow{P}0\;\text{as }N\to\infty$ (12)
Step 2: Convergence of the Level Set Estimation
We now analyze the error introduced by the Monte Carlo approximation. The high density region is the minimum set of points where the sum of the scores equals $1-\alpha$. In order to assemble the volume minimizing set, we would clearly select the points with the highest scores, leading to the thresholding condition above. Let $\xi=\hat{f}_{N}(Z)$ be the random variable representing the density score of a sample drawn from $\hat{f}_{N}$. The cumulative distribution function (CDF) of the scores is given by $G_{N}(t)=\mathbb{P}(\xi\leq t)$. We can express the thresholding condition in terms of the CDF:
$G_{N}(t_{N})=\mathbb{P}(\xi\leq t_{N})=\alpha$ (13)
Thus, the theoretical threshold $t_{N}$ is exactly the $\alpha$-quantile of the distribution of density scores.
The algorithm approximates the score CDF by drawing $M$ i.i.d. samples $z_{1},\dots,z_{M}\sim\hat{f}_{N}$ and computing their scores $\xi_{i}=\hat{f}_{N}(z_{i})$. Let $\hat{G}_{M}(t)=\frac{1}{M}\sum_{i=1}^{M}\mathbb{I}(\xi_{i}\leq t)$ be the empirical cumulative distribution function of the scores. The algorithm selects $\hat{\tau}_{M}$ as the empirical $\alpha$-quantile: $\hat{\tau}_{M}=\hat{G}_{M}^{-1}(\alpha)$, defining the estimated region $\hat{R}_{N,M}=\{x:\hat{f}_{N}(x)\geq\hat{\tau}_{M}\}$.
By the Glivenko-Cantelli theorem (Theorem B.4), the empirical CDF converges uniformly to the true CDF $G_{N}$ almost surely as $M\to\infty$. Therefore, the empirical quantile converges to the true quantile:
$|\hat{\tau}_{M}-t_{N}|\xrightarrow{a.s.}0$ (14)
Since we utilize a Gaussian kernel, the estimated density $\hat{f}_{N}$ is continuous and has no flat regions. Therefore, the measure of the symmetric difference between the level sets is continuous with respect to the threshold. The convergence of the threshold implies the convergence of the regions:
$\lambda(\hat{R}_{N,M}\Delta R_{N})\leq\lambda(\{x:|\hat{f}_{N}(x)-t_{N}|\leq|\hat{\tau}_{M}-t_{N}|\})\xrightarrow{a.s.}0$ (15)
Conclusion Combining the results, we have that as $N\to\infty$: $\hat{f}_{N}\to f\;\text{(almost surely)}$, and that as $M\to\infty$: $\hat{R}_{N,M}\to R_{N}\;\text{}$, where the set $R_{N}$ is defined by the KDE $\hat{f}_{N}$. Since $\hat{f}_{N}$ converges to $f$, the set $R_{N}$ converges to the true set $R_{\alpha}$. Therefore, we conclude that:
$\lambda(\hat{R}_{N,M}\Delta R_{\alpha})\xrightarrow{P}0$ (16)
Remark B.0.
The above proof shows that under the same conditions as Theorem B.1, the MCKDE procedure is asymptotically correct, in that it will approach the true level set as the number of samples goes to infinity.
Theorem B.3 (KDE Convergence, Theorem A, Silverman (1978)).
Assuming that the conditions in Theorem B.1 hold, the Kernel Density Estimator $\hat{f}_{N}$ converges uniformly to the true density $f$ almost surely as $N\to\infty$:
$\limsup_{N\to\infty}|\hat{f}_{N}(x)-f(x)|=0$ (17)
Theorem B.4 (Glivenko-Cantelli Theorem as written in Vaart (1998)).
@@
Corollary B.5 (Corollary 2.1 from Cadre (2006)).
This corollary is a consequence of Cadre (2006, Theorem 2.1) Let $d\geq 2$ be the dimension of the data, and $(a_{N})_{N}$ be a sequence of positive real numbers such that $a_{N}\to 0$. Let $t^{(p)}$ be the unique real number such that $\lambda_{f}(L(f,t^{(p)}))=p$, and $t_{N}^{(p)}$ be the unique real number such that $\lambda_{\hat{f}_{N}}(L(\hat{f}_{N},t_{N}^{(p)}))=p$. If the assumptions of Theorem B.1 hold, then for almost every $t\geq 0$ we have that:
$\lambda(\hat{R}_{M,N,p}\Delta R_{p})\xrightarrow{P}\frac{\sqrt{t_{N}^{(p)}}}{\sqrt{Nh^{d}}\varphi_{N}}\sqrt{\frac{2t}{\pi}\int K^{2}d\lambda}$ (19)
Where $\varphi_{N}=a_{N}/\lambda(\hat{R}_{M,N,p}-\hat{R}_{M,N,p+a_{N}})$
Note that in this corollary, the square root of the kernel term is a constant that depends on the choice of kernel, and the term $\sqrt{Nh^{d}}$ grows to infinity as $N$ increases, ensuring that the estimation error decreases with larger sample sizes. The term $t_{N}^{(p)}$ represents the density threshold corresponding to the level set, which is bounded for a well-defined density function. Thus, the corollary indicates that the estimation error diminishes as the sample size increases. The asymptotic behaviour described in Corollary B.5 is that
$\lambda(\hat{R}_{M,N,p}\Delta R_{p})\xrightarrow{P}O\left(1/\sqrt{Nh^{d}}\right)$ (20)
Appendix C Theoretical Guarantees
C.1 Regret Guarantee
We present the theorem that yields Remark 5.2
Theorem C.1 (Theorem 4, (Gibbs & Candès, 2024)).
@@ C.1 Regret Guarantee
$\begin{split}\frac{1}{|W|}\sum_{t=r}^{s}\ell((\beta_{t},\bar{\alpha}_{t}))-\frac{1}{|W|}\sum_{t=r}^{s}\ell((\beta_{t},\alpha^{*}_{t}))\leq O\left(\frac{\log(|W|)}{|W|}\right)+O\left(\right)\sqrt{\frac{\sum_{t=r+1}^{s}\left|\alpha_{t}^{*}-\alpha_{t-1}^{*}\right|+1}{|W|}},\gamma_{1}.\end{split}$ (22)
This derivation is due to Gibbs & Candès (2024)
C.2 Long Term Coverage
Theorem C.2 (Theorem 6, (Gibbs & Candès, 2024)).
Consider a modified version of Algorithm 3 in which on iteration $t$ the parameters $\eta$ and $\sigma$ are replaced by values $\eta_{t}$ and $\sigma_{t}$. Let $\gamma_{\min}:=\min_{i}\gamma_{i}$$\gamma_{\max}:=\max_{i}\gamma_{i}$, and that $\lim_{t\to\infty}\eta_{t}=\lim_{t\to\infty}\sigma_{t}=0$. Then,
@@ C.2 Long Term Coverage
C.3 Score Function Stability
See 5.3

### v2

@@ C.3 Dynamically Tuned Adaptive Conformal Inference (DtACI)
C.4 High Density Region Identification
@@
Under standard KDE regularity conditions and sufficiently dense Monte Carlo sampling, the retained high-density samples approximate the corresponding KDE level set. We therefore model the score-optimization step through Assumption C.1.
Assumption C.1 (Consistent residual-template estimation).
At score-update times $t_{m}$, let $\widehat{R}_{m}$ be the residual set returned by the score-optimization procedure and let $K_{m}=\operatorname{conv}(\widehat{R}_{m})$. Let $R_{\star}=\{z:f_{\star}(z)\geq\tau_{\star}\}$ and $K_{\star}=\operatorname{conv}(R_{\star})$. We assume
$d_{H}(K_{m},K_{\star})\xrightarrow{p}0.$
Theorem C.2 (Glivenko-Cantelli Theorem as written in Vaart [27]).
@@
C.5 Computational Complexity Analysis of AdaptNC
Theorem C.3 (Computational Complexity of AdaptNC).
AdaptNC consists of two components: a threshold update executed at every timestep and a score function update executed every $t_{s}$ timesteps. Let $W$ denote the window size, $k$ the number of experts, $M$ the number of Monte Carlo samples, $d$ the residual dimension, $r$ the number of facets of the convex hull, and $n=|\hat{R}_{t}|$ the number of retained high-density points. Then, the per-timestep cost of the threshold update is $O(r+W+k)$, and the cost of the score function update is $O\!\left(MWd+M\log M+n\log n+W(r+k)\right)$, which is dominated by the KDE evaluation term $O(MWd)$ in the low-dimensional regime considered here.
Under the assumption $d\leq 3$, the amortized per-timestep cost of AdaptNC is
$O\!\left(W+k+MWd/t_{s}\right).$
Proof.
We analyze the computational cost of AdaptNC by decomposing it into two components: the threshold update and the score function update. We assume that the residual dimensionality satisfies $d\leq 3$, which holds for all experiments and typical real-world trajectory prediction tasks.
Threshold update: The threshold update executes at every timestep and consists of three operations. First, the nonconformity score $s(X_{t},Y_{t};\theta_{t})$ is evaluated for the current observation, requiring $O(r)$ time where $r$ is the number of facets of the convex hull. Second, the inverse quantile $\beta_{t}$ is computed by evaluating the score against the rolling window $\mathcal{D}_{t}$, requiring $O(W)$ time. Third, the $k$ expert weights are updated via exponential reweighting and each expert’s quantile $\alpha_{t}^{i}$ is updated, requiring $O(k)$ time. The total per-timestep cost of the fast path is therefore $O(r+W+k)$.
Score function update: The score function update executes every $t_{s}$ timesteps and consists of three stages: reweighting, score optimization, and replay.
Reweighting: Computing the adaptive weights $\omega_{\tau}$ for all history buffer entries requires evaluating the exponential decay across $k$ experts for each of the $W$ buffered points, yielding $O(kW)$.
Score optimization: This stage comprises density estimation and shape template fitting. The weighted KDE is constructed from $W$ samples in $O(Wd)$ time. Drawing $M$ Monte Carlo samples from the KDE costs $O(Md)$. Evaluating the KDE at each Monte Carlo sample requires summing over all $W$ kernel centers, yielding $O(MWd)$. Selecting the top $1-\alpha$ fraction by density requires sorting in $O(M\log M)$. Finally, the QuickHull algorithm fits the convex hull to the $n$ retained high-density points in $O(n\log n)$ expected time. The dominant cost of this stage is the KDE evaluation at $O(MWd)$.
Replay: The replay mechanism re-initializes the DtACI expert chain and processes the $W$ most recent observations under the updated score function. Each replayed observation requires $O(r+k)$ for score evaluation and expert updates, yielding a total replay cost of $O(W(r+k))$.
The total cost of the score update is $O(MWd+M\log M+n\log n+W(r+k))$, dominated by the KDE evaluation term $O(MWd)$.
Amortized cost: Since the score function update is executed every $t_{s}$ timesteps, its per timestep cost amortizes to: $O\!\left(MWd/t_{s}\right)$. Adding the threshold update cost yields the amortized per-timestep cost of AdaptNC:
$O\!\left(W+k+MWd/t_{s}\right)$
Appendix D Theoretical Details and Proofs
D.1 Background ACI and DtACI Guarantees
We present the theorem that yields Remark˜D.2
Theorem D.1 (Theorem 4, [11]).
@@ D.1 Background ACI and DtACI Guarantees
$\begin{split}\frac{1}{|W|}\sum_{t=r}^{s}\ell((\beta_{t},\bar{\alpha}_{t}))-\frac{1}{|W|}\sum_{t=r}^{s}\ell((\beta_{t},\alpha^{*}_{t}))\leq O\left(\sqrt{\frac{\log(|W|)}{|W|}}\right)+O\left(\sqrt{\frac{\sum_{t=r+1}^{s}\left|\alpha_{t}^{*}-\alpha_{t-1}^{*}\right|+1}{|W|}}\right).\end{split}$ (10)
This derivation is due to Gibbs and Candès [11]
Theorem D.3 (Theorem 6, [11]).
Consider a modified version of Algorithm˜3 in which on iteration $t$ the parameters $\eta$ and $\sigma$ are replaced by values $\eta_{t}$ and $\sigma_{t}$. Let $\gamma_{\min}:=\min_{i}\gamma_{i}$$\gamma_{\max}:=\max_{i}\gamma_{i}$, and that $\lim_{t\to\infty}\eta_{t}=\lim_{t\to\infty}\sigma_{t}=0$. Then,
@@
D.2 HDR and Score-Stability Proofs
Assumption D.4 (Regularity for HDR stability).
Let $\mathcal{Z}_{t}$ denote the effective weighted residual distribution used by AdaptNC at score-update time $t$, and let $f_{t}$ denote its density. Let $\mathcal{Z}_{\star}$ be the limiting weighted residual distribution with density $f_{\star}$.
  1. 1.
    The residual-distribution convergence is strong enough to imply uniform density convergence:
    $\|f_{t}-f_{\star}\|_{\infty}\to 0.$
  2. 2.
    The KDE kernel, bandwidth sequence, and weighted effective sample size satisfy the standard conditions for uniform KDE consistency:
    $\|\widehat{f}_{t}-f_{t}\|_{\infty}\xrightarrow{p}0.$
  3. 3.
    The density-score quantile at level $\alpha$ is regular: if
    $G_{\star}(u):=\mathbb{P}_{Z\sim\mathcal{Z}_{\star}}(f_{\star}(Z)\leq u),$
    then $G_{\star}$ is continuous and strictly increasing in a neighborhood of $\tau_{\star}:=G_{\star}^{-1}(\alpha)$.
  4. 4.
    The limiting HDR boundary is regular: there exist constants $c>0$ and $\varepsilon_{0}>0$ such that
    $\|\nabla f_{\star}(z)\|_{2}\geq c\qquad\text{whenever }|f_{\star}(z)-\tau_{\star}|\leq\varepsilon_{0}.$
  5. 5.
    The relevant HDRs are nonempty compact subsets of $\mathbb{R}^{p}$.
  6. 6.
    There exists $L_{\mathrm{vol}}<\infty$ such that for all $x$, all thresholds $q$ in the relevant range, and all sufficiently small $\varepsilon>0$
    $\operatorname{Vol}\{y:q-\varepsilon\leq s_{\star}(x,y)\leq q+\varepsilon\}\leq L_{\mathrm{vol}}\varepsilon.$
Lemma D.5 (Hausdorff consistency of the MCKDE high-density region).
Let
$R_{\star}:=\{z\in\Omega:f_{\star}(z)\geq\tau_{\star}\}$
be the limiting $(1-\alpha)$ high-density residual region, where $\tau_{\star}$ is chosen so that
$\int_{R_{\star}}f_{\star}(z)\,dz=\int_{\Omega}\mathbf{1}\{f_{\star}(z)\geq\tau_{\star}\}f_{\star}(z)\,dz=1-\alpha.$
At score-update time $t$, let $\widehat{f}_{t}$ be the KDE formed from $N_{t}$ weighted residual samples, and let $\widehat{\tau}_{t}$ be the Monte Carlo estimate of the density threshold obtained from $M_{t}$ samples. Define the MCKDE estimated high-density region
$\widehat{R}_{t}:=\{z\in\Omega:\widehat{f}_{t}(z)\geq\widehat{\tau}_{t}\}.$
Under Assumptions 5.1 and D.4,
$d_{H}(\widehat{R}_{m},R_{\star})\xrightarrow{p}0.$